OpenStudy (anonymous):

what is the sum of the two values of x that satisfy the equation 4x^2 -3x-1=0 ?

OpenStudy (zzr0ck3r):

these are the same questions over and over and over, you need to learn to find the solutions to the quadratic equation. It will not go away:)

OpenStudy (ankitshaw):

ya its the same type of que..

OpenStudy (anonymous):

i tried and its not the right answer!

OpenStudy (ankitshaw):

what are the values of x u found ?...

OpenStudy (anonymous):

x= -1/4 & x=1

OpenStudy (ankitshaw):

its correct then what is the problem?

OpenStudy (anonymous):

The answers you must choose are a) 0.75 B) 1.25 C) 2.5 D) -0.75 E) -1.25

OpenStudy (ankitshaw):

so get the sum of the values of x and get it in decimals .. u will get the answer

OpenStudy (anonymous):

it dont work .

OpenStudy (ankitshaw):

what ans u got?

OpenStudy (anonymous):

0.25

OpenStudy (ankitshaw):

seriously .. -1/4 + 1 = 3/4 = .75